Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2016-11-25

Observers: Felipe Olivares, Lingzhi Wang, Lixin, Yu
Support team: Lluís Galbany
TAT team: Lluís Galbany, Janet Chen

The night is not photometric

Note: telescope was closed until 05:04 UT because of the weather. We couldn't take evening skyflats.

Weather conditions:
Start: moderate humidity, strong wind
Middle: strong wind, pointing restriction, MeteoMonitor stops working at 2:03 UT
End: strong wind

05:04UT Image analysis
05:37UT SN2016iae, SN2016ija follow-up, gr11 with 1.0" slit, 2x1800s, seeing ~1", Airmass 1.03
07:02UT SN2016iae, SN2016ija follow-up, gr16 with 1.0" slit, 2x1800s, seeing ~1", Airmass 1.18
08:21UT Quartz lamp for internal flat
08:28UT End of night

Daytime calibrations: 
EFOSC2:
PESSTO_bias
PESSTO_calib
Comments